Indicate how the concentration of each species in the chemical equation will change to reestablish equilibrium after a decrease in pressure or volume. Click in each box to toggle the relative change in concentration (up=increases, down-decreases, leave the box blank for no change). CH,(g) +2H,S(g) 근 CS2(g) +4H2 g) a) after decreasing the pressure b) after decreasing the volume of the reaction vessel
Answer #1


CH4(g) + 2H2S(g) <===> CS2(g) + 4H2(g)

a) after decreasing the pressure , equilibrium shifts towards more number of moles side.

in the given equation.

Total no of moles of products = 4+1 = 5

Total no of moles of reactants = 2+1 = 3

so that, equilibrium shifts towards Right.

therefore,

concentration of CS2,H2 increases and concentration of CH4,H2S decreases.

b) after decreasing the volume of the reaction vessel. the pressure inside the vessel increases.so that equilibrium shifts towards less no of moles side(reactants).

therefore,

concentration of CS2,H2 decreases and concentration of CH4,H2S increases.